if the given samples are 45, 47, 48, 51, 50, 55, 49, 48, 55, and 54. compute the point estimate of the population variance.




1
Expert's answer
2022-04-27T16:48:57-0400

"\mu=(45+47+48+51+50+55+49+48+55+54)/10=50.2"

"D=\frac{\sum(x-\mu)^2}{n-1}=\frac{27.04+10.24+4.84+0.16+0.04+23.04+1.44+4.84+23.04+14.44}{10-1}=12.1"